A 1.25 kg squirrel runs east across the ground for 7.00 meters. She climbs, 0.50 meters before running west on a limb for 3.25 meters. She jumps up .85 meters onto another branch, travels west 3.75 meters. She stops and eats a nut. Calculate net work. What is the squirrel’s gravitational potential energy?
1
Expert's answer
2020-11-10T07:04:29-0500

If we ignore friction that the squirrel experienced while traveled along the horizontal, the work is done only when it jumped up and climbed the tree. Indeed, the work is


"W=Fd\text{ cos}\theta,"

where "\theta" is the angle between the force and displacement vectors. Thus, since the only force is the force of gravity (acting downward), the work is done only on vertical paths:


"W_1=mgh_1=1.25\cdot9.8\cdot0.5=6.13\text{ J}.\\
W_2=mgh_2=1.25\cdot9.8\cdot0.85=10.4\text{ J}.\\
W_{\text{net}}=W_1+W_2=6.13+10.4=16.5\text{ J}."

The gravitational potential energy above the ground is


"GPE=mgh=mg(h_1+h_2)=16.5\text{ J}."